OpenStudy (anonymous):

Solve for b 2*log(b)/log(a)+4*log(a)/log(b) = 9 where a>b>1

OpenStudy (anonymous):

Actually \[2\log_{a}b+4\log_{b}a = 9\]

hartnn (hartnn):

use, \(\large\log_xy=\dfrac{1}{\log_yx}\) then put \(\log_ba=x\) you'll get a quadratic in 'x', which you can solve easily.

hartnn (hartnn):

see if you get \(4x^2-9x+2=0\)

OpenStudy (anonymous):

is "a" a constant..or we have to find some numerical value of "a" also..otherwise the ans will be in terms of "a"

hartnn (hartnn):

^yes, the ans will be in terms of "a"

OpenStudy (anonymous):

that means we dont have to find the value of "a"..can we find some numerical value of "a" by any means??

hartnn (hartnn):

2 unknows and only 1 equation. i don't think it'll be possible to find numerical value of a or b

OpenStudy (anonymous):

yeah,i also think so..